# Channel Capacity of Near-Field Multiuser Communications

### v1

@@ Abstract
The channel capacity of near-field (NF) communications is characterized by considering three types of multiuser channels: i) multiple access channel (MAC), ii) broadcast channel (BC), and iii) multicast channel (MC). For NF MAC and BC, closed-form expressions are derived for the sum-rate capacity as well as the capacity region under a two-user scenario. These results are further extended to scenarios with an arbitrary number of users. For NF MC, closed-form expressions are derived for the two-user channel capacity and the capacity upper bound with more users. Further insights are gleaned by exploring special cases, including scenarios with infinitely large array apertures, co-directional users, and linear arrays. Theoretical and numerical results are presented and compared with far-field communications to demonstrate that: i) the NF capacity of these three channels converges to finite values rather than growing unboundedly as the number of array elements increases; ii) the capacity of the MAC and BC with co-directional users can be improved by using the additional range dimensions in NF channels to reduce inter-user interference (IUI); and iii) the MC capacity benefits less from the NF effect compared to the MAC and BC, as multicasting is less sensitive to IUI.
@@ I Introduction
In light of recent developments in wireless networks, emerging technical trends, such as the application of extremely large-scale antenna arrays and tremendously high frequencies, significantly expand the near-field (NF) region, even to hundreds of meters [1]. It is important to emphasize that electromagnetic (EM) waves exhibit distinct propagation characteristics in the NF region compared to the far field where EM waves can be adequately approximated as planar waves. In the near field, a more precise spherical wave-based model becomes necessary [2]. Therefore, it is imperative to reevaluate the performance of multiuser systems from an NF perspective, which ensures that modeling and analysis accurately reflect these distinct propagation characteristics.
By leveraging the additional range dimensions introduced by spherical wave propagation [1], near-field communications (NFC) can manage inter-user interference (IUI) more flexibly. This capability has inspired a considerable amount of research focused on multiuser NF beamforming design; see [3, 4, 5] and the references therein for more details. On the other hand, the fundamental performance limits of NF multiuser communications have not been adequately studied. Only the achievable rates for an uplink NF multiuser channel were studied in [6], where the maximum-ratio transmission, zero-forcing, and minimum mean-squared error beamforming strategies are considered. By now, one of the most fundamental problems in NF multiuser communications remains unsolved: channel capacity characterization. This issue has long been of significant value and interest within the realm of multiuser multiple-antenna communications [7].
Research on multiuser communications has focused on several primary models that play fundamental roles in shaping the theoretical and practical landscapes of communication networks [8, 9]. Among these models, three classical ones are outlined here, which are also the main focus of this article: the multiple access channel (MAC) [10], where multiple transmitters communicate with a single receiver; the broadcast channel (BC) [11], where a single transmitter broadcasts different messages to multiple receivers; and the multicast channel (MC) [12], where a single transmitter sends a common message to multiple receivers. The MAC addresses the challenges of uplink communications, while the BC and MC capture the essence of downlink communications. Prior studies have extensively characterized the (sum-rate) capacity and capacity regions of these channels under various conditions, which unveil the nature of multiuser communications and its capacity limits; see [7, 8, 9] for further details.
@@
Motivated by existing research gaps, this article analyzes the channel capacity of NF multiuser communications in terms of the aforementioned three fundamental channels: MAC, BC, and MC. The main contributions are summarized as follows.
@@
  • Building upon the multiuser NFC framework, we derive closed-form expressions for the sum-rate channel capacities of NF MAC and BC under a two-user scenario, along with the corresponding capacity region. These results are then extended to general scenarios with an arbitrary number of users. For NF MC, we propose an optimal linear beamforming design under a two-user scenario and derive the corresponding multicast capacity. For scenarios with more users, we provide a closed-form expression for the upper bound of the NF multicast capacity.
@@
The remainder of this article is organized as follows. Section II presents the NF channel model and defines the capacity of the three multiuser channels. Then, Sections III, IV and V analyzes the NF channel capacity of MAC, BC and MC, respectively. Section VI provides numerical results to validate the derived theoretical insights. Finally, Section VII concludes the article.
@@ II System Model
Consider a narrowband single-cell multiuser system where one base station (BS) simultaneously serves a set of $K$ user terminals (UTs), as depicted in Figure 1. Each UT $k\in{\mathcal{K}}\triangleq\{1,\ldots,K\}$ is a single-antenna device, while the BS is equipped with a large-aperture uniform planar array (UPA) containing $M\gg K$ antennas. The deployment of this massive array extends the NF region, which resides all the UTs within the near field. Since NF channels are sparsely-scattered and dominated by line-of-sight (LoS) propagation [1], we consider pure-LoS propagation scenarios for a theoretical exploration of fundamental capacity limits.
@@
Regarding each UT $k\in\mathcal{K}$, they are all assumed to be equipped with a single hypothetical isotropic array element to receive or transmit signals. Let $r_{k}$ denote the propagation distance from the center of the antenna array to UT $k$, and $\theta_{k}\in[0,\pi]$ and $\phi_{k}\in[0,\pi]$ denote the associated azimuth and elevation angles, respectively. Thus, the location of UT $k$ can be expressed as ${\mathbf{r}}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Omega_{k}]^{\mathsf{T}}$, where $\Phi_{k}\triangleq\sin{\phi_{k}}\cos{\theta_{k}}$, $\Psi_{k}\triangleq\sin\phi_{k}\sin\theta_{k}$, and $\Omega_{k}\triangleq\cos{\phi_{k}}$. In particular, the distance between UT $k$ and the center of the $(m_{x},m_{z})$th array element is given by
@@ II-A Channel Model
$h_{m_{x},m_{z},k}\approx\sqrt{\frac{Ar_{k}\Psi_{k}}{4\pi r_{m_{x},m_{z},k}^{3}}}\mathrm{e}^{-\mathrm{j}\frac{2\pi}{\lambda}r_{m_{x},m_{z},k}}.$ (5)
@@
$h_{m_{x},m_{z},k}\approx\sqrt{\frac{A\Psi_{k}}{4\pi r_{k}^{2}}}\mathrm{e}^{-\mathrm{j}\frac{2\pi r_{k}}{\lambda}(1-m_{x}\epsilon_{k}\Phi_{k}-m_{z}\epsilon_{k}\Omega_{k})}.$ (6)
Comparing (5) with (6) leads to the following observation.
Remark 1.
In contrast to FF channels, spherical wavefronts introduce additional range dimensions $\{r_{m_{x},m_{z},k}\}_{m_{x},m_{z}}$ to NF channels.
@@ II-B1 MAC
In the sequel, we will analyze the NF capacity of the above three channels, and compare them with their FF counterpart to gather insights. For clarity, the following analysis will focus on the two-user scenario, while certain findings can be extended to cases involving more than two UTs.
@@ III Multiple Access Channel
In this section, we analyze the NF channel capacity of MAC by deriving its sum-rate capacity and capacity region.
III-A Sum-Rate Capacity
Given the power budget $P_{k}$ for UT $k$, the sum-rate capacity of the MAC can be written as follows:
${\mathsf{C}}_{\mathsf{MAC}}=\max_{0\leq p_{k}\leq P_{k}}\log_{2}\det\left({\mathbf{I}}+\frac{1}{\sigma^{2}}\sum\nolimits_{k=1}^{K}{p_{k}}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\right),$ (9)
which can be achieved by using point-to-point Gaussian random coding along with successive interference cancellation (SIC) decoding in some message decoding order [7]. From (9), it is readily shown that the sum-rate MAC capacity is attained when each UT transmits at the maximum power, i.e., $p_{k}=P_{k}$ for $k\in{\mathcal{K}}$, which yields
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}\det\left({\mathbf{I}}+\frac{1}{\sigma^{2}}\sum\nolimits_{k=1}^{K}{P_{k}}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\right).$ (10)
In the sequel, we derive a closed-form expression for ${\mathsf{C}}_{\mathsf{MAC}}$ by considering a two-user scenario, i.e., $K=2$, and the results will be further extended to the case of $K>2$ in Section III-D. By defining $\rho\triangleq\frac{\lvert{\mathbf{h}}_{1}^{\mathsf{H}}{\mathbf{h}}_{2}\rvert^{2}}{\lVert{\mathbf{h}}_{1}\rVert^{2}\lVert{\mathbf{h}}_{2}\rVert^{2}}\in[0,1]$ as the channel correlation factor (CCF), we obtain the following lemma.
@@ III-A Sum-Rate Capacity
The sum-rate capacity of the two-user MAC can be expressed as follows:
$\begin{split}\mathsf{C}_{\mathsf{MAC}}&=\log_{2}(1+{\sigma^{-2}}({P_{1}\|\mathbf{h}_{1}\|^{2}+P_{2}\|\mathbf{h}_{2}\|^{2}})\\ &~{}~{}+{\sigma^{-4}}{P_{1}P_{2}\|\mathbf{h}_{1}\|^{2}\|\mathbf{h}_{2}\|^{2}}(1-\rho)).\end{split}$ (11)
@@
Remark 2.
The results in (11) suggest that the sum-rate capacity is influenced by the CCF between the two UTs, which reflects the impact of IUI. An ideal scenario occurs when the channels for the two UTs are orthogonal, which yields $\rho=0$ and indicates no IUI. This establishes an upper bound for $\mathsf{C}_{\mathsf{MAC}}$, which can be expressed as follows:
$\mathsf{C}_{\mathsf{MAC}}$ $\leq\log_{2}(1+{\sigma^{-2}}({P_{1}\|\mathbf{h}_{1}\|^{2}+P_{2}\|\mathbf{h}_{2}\|^{2}})$
$~{}~{}+{\sigma^{-4}}{P_{1}P_{2}\|\mathbf{h}_{1}\|^{2}\|\mathbf{h}_{2}\|^{2}})$
$=\sum\nolimits_{k=1}^{2}\log_{2}(1+{\sigma^{-2}}P_{k}\|\mathbf{h}_{k}\|^{2}).$ (12)
By further incorporating the NF channel model (5) into the analysis, we derive the NF MAC capacity as follows.
@@
The sum-rate capacity of the MAC under the NF model is given by
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}\!\left(1\!+\frac{P_{1}G_{1}\!+\!P_{2}G_{2}}{\sigma^{2}}+\frac{P_{1}P_{2}G_{1}G_{2}\bar{\rho}}{\sigma^{4}}\right),$ (13)
where $\bar{\rho}=1-\rho_{{}_{\mathsf{N}}}$. Specifically,
$G_{k}=\frac{\xi}{4\pi}\sum\nolimits_{x\in\{\frac{M_{x}\epsilon_{k}}{2}\pm\Phi_{k}\}}{\sum\nolimits_{z\in\{\frac{M_{z}\epsilon_{k}}{2}\pm\Omega_{k}\}}{\delta_{k}\left(x,z\right)}}$ (14)
denotes the channel gain of UT $k=1,2$. Furthermore,
$\begin{split}\rho_{{}_{\mathsf{N}}}&=\prod\nolimits_{k=1}^{2}{\frac{d^{2}r_{k}\Psi_{k}}{\sum_{x\in\mathcal{X}_{k}}{\sum_{z\in\mathcal{Z}_{k}}{\delta_{k}\left(x,z\right)}}}}\\ &\times\left[\left(\sum\nolimits_{m_{x}\in\mathcal{M}_{x}}{\sum\nolimits_{m_{z}\in\mathcal{M}_{z}}{\Upsilon_{m_{x},m_{z}}^{-\frac{3}{2}}\cos\Delta_{m_{x},m_{z}}}}\right)^{2}\right.\\ &\left.+\left(\sum\nolimits_{m_{x}\in\mathcal{M}_{x}}{\sum\nolimits_{m_{z}\in\mathcal{M}_{z}}{\Upsilon_{m_{x},m_{z}}^{-\frac{3}{2}}\sin\Delta_{m_{x},m_{z}}}}\right)^{2}\right]\end{split}$ (15)
denotes the CCF under the NF model, where $\Upsilon_{m_{x},m_{z}}=r_{m_{x},m_{z},1}r_{m_{x},m_{z},2}$, and $\Delta_{m_{x},m_{z}}=\frac{2\pi}{\lambda}\left(r_{m_{x},m_{z},1}-r_{m_{x},m_{z},2}\right)$.
@@
We next aim to gain further insights into $\mathsf{C}_{\mathsf{MAC}}$ by considering an infinitely large array aperture. Specifically, when $M_{x},M_{z}\rightarrow\infty$, we have
$\lim_{M_{x},M_{z}\rightarrow\infty}G_{k}=\frac{\xi}{4\pi}4\times\lim_{x,z\rightarrow\infty}\delta_{k}\left(x,z\right)=\frac{\xi}{2},$ (16)
where $\xi=\frac{A}{d^{2}}\in\left(0,1\right]$ denotes the array occupation ratio (AOR) that measures the proportion of the entire UPA area occupied by antennas. Combining (12) with (16) gives
$\lim_{M_{x},M_{z}\rightarrow\infty}\mathsf{C}_{\mathsf{MAC}}\leq\sum\nolimits_{k=1}^{2}\log_{2}(1+P_{k}\xi/(2\sigma^{2})),$ (17)
which suggests that the NF MAC capacity is capped as $M_{x},M_{z}\rightarrow\infty$.
In contrast to $\lim_{M_{x},{M}_{z}\rightarrow\infty}G_{k}$, $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{{}_{\mathsf{N}}}$ is computationally intractable. However, numerical results presented in [19] demonstrate that $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{{}_{\mathsf{N}}}\ll 1$, which is also verified by the results in Section VI. This observation suggests that as the array aperture increases, NF channels of UTs positioned at different locations tend to become orthogonal, which presents an opportunity to mitigate IUI.
@@
$\lim_{M_{x},M_{z}\rightarrow\infty}\mathsf{C}_{\mathsf{MAC}}\approx\sum\nolimits_{k=1}^{2}\log_{2}(1+P_{k}\xi/(2\sigma^{2})).$ (18)
@@
Equation (18) is derived using (16) and the fact that $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{{}_{\mathsf{N}}}\ll 1$. ∎
@@
We then consider a special case where the BS uses a uniform linear array (ULA), i.e., $M_{x}=1$ and $M_{z}=M$.
@@
When using a ULA, the channel gains satisfy
$G_{k}=\frac{\xi\epsilon_{k}\sin\phi_{k}\Xi}{4\pi\sin\theta_{k}},$ (19)
@@
$\lim_{M\rightarrow\infty}\!\mathsf{C}_{\mathsf{MAC}}\!\approx\log_{2}$ $\!\left(\!1\!+\!\frac{P_{1}\epsilon_{1}\sin\phi_{1}\sin\theta_{2}\!+\!P_{2}\epsilon_{2}\sin\phi_{2}\sin\theta_{1}}{2\pi\sigma^{2}\sin\theta_{1}\sin\theta_{2}}\xi\right.$
$\left.+\frac{P_{1}P_{2}\epsilon_{1}\epsilon_{2}\sin\phi_{1}\sin\phi_{2}\xi^{2}}{4\pi^{2}\sigma^{4}\sin\theta_{1}\sin\theta_{2}}\right),$ (20)

### v2

# Channel Capacity of Near-Field Line-of-Sight Multiuser Communications

@@ Abstract
The channel capacity of near-field (NF) communications is characterized by considering three types of line-of-sight multiuser channels: i) multiple access channel (MAC), ii) broadcast channel (BC), and iii) multicast channel (MC). For NF MAC and BC, closed-form expressions are derived for the sum-rate capacity as well as the capacity region under a two-user scenario. These results are further extended to scenarios with an arbitrary number of users. For NF MC, closed-form expressions are derived for the two-user channel capacity and the capacity upper bound with more users. Further insights are gleaned by exploring special cases, including scenarios with infinitely large array apertures, co-directional users, and linear arrays. For comparison, the MAC and BC sum-rates achieved by typical linear combiners and precoders are also analyzed. Theoretical and numerical results are presented and compared with far-field communications to demonstrate that: i) the NF capacity of these three channels converges to finite values rather than growing unboundedly as the number of array elements increases; ii) the capacity of the MAC and BC with co-directional users can be improved by using the additional range dimensions in NF channels to reduce inter-user interference (IUI); and iii) the MC capacity benefits less from the NF effect compared to the MAC and BC, as multicasting is less sensitive to IUI.
@@ I Introduction
In light of recent developments in wireless networks, emerging technical trends, such as the application of extremely large-scale antenna arrays and tremendously high frequencies, significantly expand the near-field (NF) region, even to hundreds of meters [2, 3]. It is important to emphasize that electromagnetic (EM) waves exhibit distinct propagation characteristics in the NF region compared to the far field where EM waves can be adequately approximated as planar waves. In the near field, a more precise spherical wave-based model becomes necessary [4]. Therefore, it is imperative to reevaluate the performance of multiuser systems from an NF perspective, which ensures that modeling and analysis accurately reflect these distinct propagation characteristics.
By leveraging the additional range dimensions introduced by spherical wave propagation [2], near-field communications (NFC) can manage inter-user interference (IUI) more flexibly. This capability has inspired a considerable amount of research focused on multiuser NF beamforming design; see [5, 6, 7, 8, 9] for more details. On the other hand, the fundamental performance limits of NF multiuser communications have not been adequately studied. Only the achievable rates for an uplink NF multiuser channel were studied in [10], where the maximum-ratio transmission, zero-forcing, and minimum mean-squared error beamforming strategies are considered. By now, one of the most fundamental problems in NF multiuser communications remains unsolved: channel capacity characterization. This issue has long been of significant value and interest within the realm of multiuser multiple-antenna communications [11].
Research on multiuser communications has focused on several primary models that play fundamental roles in shaping the theoretical and practical landscapes of communication networks [12, 13]. Among these models, three classical ones are outlined here, which are also the main focus of this article: the multiple access channel (MAC) [14, 15], where multiple transmitters communicate with a single receiver; the broadcast channel (BC) [16, 17], where a single transmitter broadcasts different messages to multiple receivers; and the multicast channel (MC) [18, 19], where a single transmitter sends a common message to multiple receivers. The MAC addresses the challenges of uplink communications, while the BC and MC capture the essence of downlink communications. Prior studies have extensively characterized the (sum-rate) capacity and capacity regions of these channels under various conditions, which unveil the nature of multiuser communications and its capacity limits; see [11, 12, 13] and the references therein for further details.
@@
Motivated by existing research gaps, this article analyzes the channel capacity of NF multiuser communications in terms of the aforementioned three fundamental channels: MAC, BC, and MC [^1] It is worth noting that the capacity regions of the MAC and BC are formally equivalent to those classically studied in information theory. The main contributions of this paper lie in revisiting these results within the NF setup, which allows us to explore the impact of NF effects on channel capacity. More specifically, while the basic expressions for the capacity of these channels are well known and can be found in many textbooks, their specific forms under NF channel models have not been presented in the current literature. Additionally, a detailed discussion of the effects of NF and spherical-wave propagation on channel capacity is still lacking. This gap in the literature has motivated our work.. To facilitate theoretical investigations into fundamental performance limits and asymptotic behaviors, this discussion is focused on line-of-sight (LoS) channels. Extensions to non-LoS NF channels will be addressed in future work. The main contributions are summarized as follows.
@@
  • Building upon the multiuser NFC framework, we derive closed-form expressions for the sum-rate channel capacity of NF MAC and BC under a two-user scenario, along with the corresponding capacity region. These results are then extended to more general scenarios with an arbitrary number of users. For comparison, we also analyze the MAC and BC sum-rates achieved by typical linear combiners and precoders. In the context of NF MC, we propose an optimal linear beamforming design for a two-user scenario and derive the corresponding multicast capacity. For scenarios with more users, we provide a closed-form expression for the upper bound of the NF multicast capacity.
@@
The remainder of this article is organized as follows. Section II presents the NF channel model and defines the capacity of the three multiuser channels. Then, Sections III, IV and V analyzes the NF channel capacity of MAC, BC and MC, respectively. Section VI provides numerical results to validate the derived insights. Finally, Section VII concludes the article.
@@ II System Model
Consider a narrowband single-cell multiuser system where one base station (BS) simultaneously serves a set of $K$ user terminals (UTs), as depicted in Figure 1. Each UT $k\in{\mathcal{K}}\triangleq\{1,\ldots,K\}$ is a single-antenna device, while the BS is equipped with a large-aperture uniform planar array (UPA) containing $M\gg K$ antennas. The deployment of this massive array extends the NF region, which resides all the UTs within the near field. Since NF channels are sparsely-scattered and dominated by LoS propagation [2], we consider pure-LoS propagation scenarios for a theoretical exploration of fundamental capacity limits.
@@
Regarding each UT $k\in\mathcal{K}$, they are all assumed to be equipped with a single hypothetical isotropic array element to receive or transmit signals. Let $r_{k}$ denote the propagation distance from the center of the antenna array to UT $k$, and $\theta_{k}\in(0,\pi)$ and $\phi_{k}\in(0,\pi)$ denote the associated azimuth and elevation angles, respectively. Thus, the location of UT $k$ can be expressed as ${\mathbf{r}}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Omega_{k}]^{\mathsf{T}}$, where $\Phi_{k}\triangleq\sin{\phi_{k}}\cos{\theta_{k}}$, $\Psi_{k}\triangleq\sin\phi_{k}\sin\theta_{k}$, and $\Omega_{k}\triangleq\cos{\phi_{k}}$. In particular, the distance between UT $k$ and the center of the $(m_{x},m_{z})$th array element is given by
@@ II-A Channel Model
$h_{m_{x},m_{z},k}\approx\sqrt{{Ar_{k}\Psi_{k}}/{(4\pi r_{m_{x},m_{z},k}^{3})}}\mathrm{e}^{-\mathrm{j}\frac{2\pi}{\lambda}r_{m_{x},m_{z},k}}.$ (5)
@@
$h_{m_{x},m_{z},k}\approx\sqrt{{A\Psi_{k}}/({4\pi r_{k}^{2}})}\mathrm{e}^{-\mathrm{j}\frac{2\pi r_{k}(1-m_{x}\epsilon_{k}\Phi_{k}-m_{z}\epsilon_{k}\Omega_{k})}{\lambda}}.$ (6)
By comparing (5) with (6), we can observe that, in contrast to FF channels, spherical wavefronts introduce additional range dimensions $\{r_{m_{x},m_{z},k}\}_{m_{x},m_{z}}$ to NF channels.
@@ II-B1 MAC
In the sequel, we will analyze the NF capacity of the above three channels, and compare them with their FF counterpart to gather insights.
@@ III Multiple Access Channel
In this section, we analyze the NF channel capacity of a MAC by deriving its sum-rate capacity and capacity region. The sum-rate capacity of the MAC in (7) is given by [27]
${\mathsf{C}}_{\mathsf{MAC}}=\max_{0\leq p_{k}\leq P_{k}}\log_{2}\det\Big{(}{\mathbf{I}}_{M}+\sum\nolimits_{k=1}^{K}\frac{p_{k}}{\sigma^{2}}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\Big{)},$ (9)
where $P_{k}$ represents the power budget of UT $k$. This capacity is achieved using point-to-point Gaussian random coding along with successive interference cancellation (SIC) decoding in a specific message decoding order. It is evident that the MAC sum-rate capacity is maximized when $p_{k}=P_{k}$ for $k\in{\mathcal{K}}$. By defining $\gamma_{k}\triangleq\frac{P_{k}}{\sigma^{2}}$ as the transmit signal-to-noise ratio (SNR), we can rewrite the sum-rate capacity as follows:
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}\det\Big{(}{\mathbf{I}}_{M}+\sum\nolimits_{k=1}^{K}\gamma_{k}{\mathbf{h}}_{k}{\mathbf{h}}_{k}^{\mathsf{H}}\Big{)}.$ (10)
Deriving an analytically tractable expression for ${\mathsf{C}}_{\mathsf{MAC}}$ under the NF model is a challenging task. As a compromise, we will focus on the two-user case to gain more insights, which will then be extended to scenarios with more than two UTs.
III-A Near-Field Capacity of the Two-User Case
III-A1 Sum-Rate Capacity
By denoting ${\mathsf{g}}_{k}=\lVert{\mathbf{h}}_{k}\rVert^{2}$ as the channel gain of UT $k$ and defining $\rho\triangleq\frac{\lvert{\mathbf{h}}_{1}^{\mathsf{H}}{\mathbf{h}}_{2}\rvert^{2}}{\lVert{\mathbf{h}}_{1}\rVert^{2}\lVert{\mathbf{h}}_{2}\rVert^{2}}\in[0,1]$ as the channel correlation factor (CCF) [^2] The CCF is an important metric that quantifies the correlation between the channels of two UTs. When $\rho=0$, the spatial channels for the UTs are orthogonal; when $\rho=1$, the spatial channels are parallel. between the two UTs, the rates of the UTs achieved by SIC with different decoding orders are calculated as follows.
@@ III-A1 Sum-Rate Capacity
When the SIC decoding order $1\rightarrow 2$ is adopted, the rates of the UTs are, respectively, given by
${\mathsf{R}}_{1}^{1\rightarrow 2}=\log_{2}\left(1+\frac{\gamma_{1}{\mathsf{g}}_{1}+\gamma_{1}\gamma_{2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}(1-\rho)}{1+\gamma_{2}{\mathsf{g}}_{2}}\right),$ (11a)
${\mathsf{R}}_{2}^{1\rightarrow 2}=\log_{2}\left(1+\gamma_{2}{\mathsf{g}}_{2}\right).$ (11b)
For the decoding order $2\rightarrow 1$, we have
${\mathsf{R}}_{1}^{2\rightarrow 1}=\log_{2}\left(1+\gamma_{1}{\mathsf{g}}_{1}\right),$ (12a)
${\mathsf{R}}_{2}^{2\rightarrow 1}=\log_{2}\left(1+\frac{\gamma_{2}{\mathsf{g}}_{2}+\gamma_{1}\gamma_{2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}(1-\rho)}{1+\gamma_{1}{\mathsf{g}}_{1}}\right).$ (12b)
@@
The sum-rate capacity of the two-user MAC is given by
$\mathsf{C}_{\mathsf{MAC}}$ $={\mathsf{R}}_{1}^{1\rightarrow 2}+{\mathsf{R}}_{2}^{1\rightarrow 2}={\mathsf{R}}_{1}^{2\rightarrow 1}+{\mathsf{R}}_{2}^{2\rightarrow 1}$
$=\log_{2}(1+{\gamma_{1}{\mathsf{g}}_{1}+\gamma_{2}{\mathsf{g}}_{2}}+{\gamma_{1}\gamma_{2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}(1-{\rho})}),$ (13)
which can also be obtained by directly simplifying (10) [28, Eq. (14)]. The results suggest that the sum-rate capacity is determined by the channel gains and the CCF. By incorporating the NF model (5) into the analysis, we derive a new expression for NF sum-rate capacity of the MAC as follows.
@@
The sum-rate capacity of the MAC under the NF model can be expressed as follows:
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}(1+\gamma_{1}{{\mathsf{g}}^{\mathsf{n}}_{1}}+\gamma_{2}{{\mathsf{g}}}^{\mathsf{n}}_{2}+{\gamma_{1}\gamma_{2}{{\mathsf{g}}^{\mathsf{n}}_{1}}{{\mathsf{g}}^{\mathsf{n}}_{2}}(1-{{\rho}_{\mathsf{n}}})}).$ (14)
For the terms appearing in (14),
$\begin{split}{{\mathsf{g}}}^{\mathsf{n}}_{k}=\frac{\xi}{4\pi}\sum_{x\in{\mathcal{X}}_{k}}\sum_{z\in{\mathcal{Z}}_{k}}\arctan\bigg{(}\frac{xz/\Psi_{k}}{\sqrt{\Psi_{k}^{2}+x^{2}+z^{2}}}\bigg{)}\end{split}$ (15)
is the NF channel gain of UT $k$, where ${\mathcal{X}}_{k}=\{\frac{M_{x}}{2}\epsilon_{k}\pm\Phi_{k}\}$, ${\mathcal{Z}}_{k}=\{\frac{M_{z}}{2}\epsilon_{k}\pm\Psi_{k}\}$, and $\xi=\frac{A}{d^{2}}\in\left(0,1\right]$ denotes the array occupation ratio (AOR). Furthermore, the CCF under the NF model can be calculated as
${\rho}_{\mathsf{n}}$ $=\prod\nolimits_{k=1}^{2}{\frac{\pi MA\Psi_{k}}{4r_{k}^{2}\overline{\mathsf{g}}_{k}T^{2}}}\left|\sum\nolimits_{t=1}^{T}\sum\nolimits_{t^{\prime}=1}^{T}{\sqrt{\left(1-\delta_{t}^{2}\right)\left(1-\delta_{t^{\prime}}^{2}\right)}}\right.$
×$\left.\times f_{1}\left(M_{x}\epsilon_{1}\delta_{t},M_{z}\epsilon_{1}\delta_{t^{\prime}}\right)f_{2}\left(M_{x}\epsilon_{1}\delta_{t},M_{z}\epsilon_{1}\delta_{t^{\prime}}\right)\right|^{2},$ (16)
where $T$ is a complexity-vs-accuracy tradeoff parameter, $\delta_{t}=\cos\left(\frac{2t-1}{2T}\pi\right)$ for $t\in\left\{1,\ldots,T\right\}$, and
$f_{1}\left(x,z\right)$ $\triangleq\frac{\mathrm{e}^{\mathrm{j}\frac{2\pi}{\lambda}r_{1}\left(x^{2}+z^{2}-2\Phi_{1}x-2\Omega_{1}z+1\right)^{\frac{1}{2}}}}{\left(x^{2}+z^{2}-2\Phi_{1}x-2\Omega_{1}z+1\right)^{\frac{3}{2}}},$ (17a)
$f_{2}\left(x,z\right)$ $\triangleq\frac{\mathrm{e}^{-\mathrm{j}\frac{2\pi}{\lambda}r_{2}\left(\upsilon^{2}x^{2}+\upsilon^{2}z^{2}-2\upsilon\Phi_{2}x-2\upsilon\Omega_{2}z+1\right)^{\frac{1}{2}}}}{\left(\upsilon^{2}x^{2}+\upsilon^{2}z^{2}-2\upsilon\Phi_{2}x-2\upsilon\Omega_{2}z+1\right)^{\frac{3}{2}}}$ (17b)
with $\upsilon={r_{1}}/{r_{2}}$.
@@
We note the CCF reflects the level of IUI, and thus an ideal scenario occurs when $\rho=0$, indicating the absence of IUI. This establishes an upper bound for $\mathsf{C}_{\mathsf{MAC}}$, which can be expressed as follows:
$\mathsf{C}_{\mathsf{MAC}}$ $\leq\log_{2}(1+{\gamma_{1}\mathsf{g}_{1}^{\mathsf{n}}+\gamma_{2}\mathsf{g}_{2}^{\mathsf{n}}}+{\gamma_{1}\gamma_{2}\mathsf{g}_{1}^{\mathsf{n}}\mathsf{g}_{2}^{\mathsf{n}}})$ (18a)
$=\sum\nolimits_{k=1}^{2}\log_{2}(1+\gamma_{k}\mathsf{g}_{k}^{\mathsf{n}}).$ (18b)
Next, we aim to gain further insights into NF MAC capacity by considering an infinitely large array aperture. Specifically, when $M_{x},M_{z}\rightarrow\infty$, we have
$\lim_{M_{x},M_{z}\rightarrow\infty}\mathsf{g}_{k}^{\mathsf{n}}$ $=\frac{4\xi}{4\pi}\lim_{x,z\rightarrow\infty}\arctan\bigg{(}\frac{xz/\Psi_{k}}{\sqrt{\Psi_{k}^{2}+x^{2}+z^{2}}}\bigg{)}$
$=\frac{4\xi}{4\pi}\frac{\pi}{2}=\frac{\xi}{2},$ (19)
which, together with (18b), yields
$\lim\nolimits_{M_{x},M_{z}\rightarrow\infty}\mathsf{C}_{\mathsf{MAC}}\leq\sum\nolimits_{k=1}^{2}\log_{2}(1+\xi\gamma_{k}/2),$ (20)
The results suggest that the NF MAC capacity is upeer bounded by a finite value as $M_{x},M_{z}\rightarrow\infty$.
In contrast to $\lim_{M_{x},{M}_{z}\rightarrow\infty}\mathsf{g}_{k}^{\mathsf{n}}$, $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{\mathsf{n}}$ is computationally intractable. However, numerical results presented in [29] demonstrate that $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{\mathsf{n}}\ll 1$, which is also verified by the results in Section VI. This observation suggests that as the array aperture increases, NF channels of UTs positioned at different locations tend to become orthogonal, which presents an opportunity to mitigate IUI.
@@
$\lim\nolimits_{M_{x},M_{z}\rightarrow\infty}\mathsf{C}_{\mathsf{MAC}}\approx\sum\nolimits_{k=1}^{2}\log_{2}(1+\xi\gamma_{k}/2).$ (21)
@@
Equation (21) is derived using (19) and the fact that $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{\mathsf{n}}\ll 1$. ∎
@@
Figure 2: Illustration of the ULA.
Turn to a special case where the BS uses a uniform linear array (ULA), i.e., $M_{x}=1$ and $M_{z}=M$, as shown in Fig 2.
@@
When using a ULA, the NF channel gains satisfy
$\mathsf{g}_{k}^{\mathsf{n}}=\frac{\xi\epsilon_{k}\sin\phi_{k}\Xi}{4\pi\sin\theta_{k}},$ (22)
@@
$\lim\nolimits_{M\rightarrow\infty}\mathsf{C}_{\mathsf{MAC}}\approx\sum\nolimits_{k=1}^{2}{\log_{2}\left(1+\frac{\gamma_{k}A\sin\phi_{k}\!}{2\pi dr_{k}\sin\theta_{k}}\right)},$ (23)